A 25 kg block is suspended by two ropes. Each rope makes an angle of 53 degrees to the horizontal. What is the magnitude of the tension force in each rope?
1
Expert's answer
2021-10-15T10:22:30-0400

Explanations & Calculations


  • According to the symmetry of the arrangement, the tension force on each rope should be equal.
  • Take it to be "\small T", and the net force along a verticle direction should be zero as the system remains in equilibrium under the two tension forces.
  • Then,

"\qquad\qquad
\begin{aligned}
\small T \sin \theta +T \sin\theta &=\small mg\\
\small 2T \sin 53 &=\small 25 kg\times 9.81 ms^{-2}\\
\small T &=\small \frac{25\times 9.81}{2\times0.799}\\
&=\small \bold{154.0 \,N}
\end{aligned}"